# Rethinking Fine-Tuning when Scaling Test-Time Compute: Limiting Confidence Improves Mathematical Reasoning

### v2

@@ Appendix C Additional results
C.2 Addition results with Llama-3-70B-base
@@
Figure 10: DCO improves on CE for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier with Llama-3-70B-base. Same as Figure 4 but with Llama-3-70B-base model. We fine-tune Llama-3-70B base models on the MATH dataset to produce direct answers. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. We plot pass@N test coverage as a function of $N$, with each curve (solid red or faint blue-green) corresponding to one fine-tuned model. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
C.3 Out-of-distribution evaluation
@@
Figure 12: DCO improves on CE for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier on out-of-distribution test set. Same as Figure 4 but evaluated on AIME24. We fine-tune Llama-3-8B base models on the MATH dataset to produce direct answers and evaluate on AIME24. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. We plot pass@N test coverage as a function of $N$, with each curve (solid red or faint blue-green) corresponding to one fine-tuned model. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
C.4 Theorem proving
In Table 6, we show additional results for model performance under the DCOSTEP objective. We find that the optimal $N_{\text{eff}}$ grows with increasing passes, agreeing with results in Sections 5.2 and 5.4. We also conduct expert iteration (Anthony et al., 2017; Polu et al., 2023) on Mathlib with theorems that do not have proof traces in the training set. We use pass@1k to prove those theorems. We find that our algorithm achieves a stronger improvement over the baseline for pass@4k after the 1st iteration. This improvement might result from the fact that the models can prove more easy theorems where the model has a higher confidence. As a result, we believe our method will perform better with expert iteration.
@@ C.4 Theorem proving
C.5 Plot of the upper bound and the lower bound
@@
C.6 Empirical evaluation of the approximately well calibrated assumption
@@
C.7 The data dependency of confidence and factor $F$
Figure 15: Model is more confident on easy problems; DCO improves test-time scaling by regularizing the easy examples. We fine-tune a Llama-3-8B-base model on the MATH dataset with CE loss and plot the model confidence $p(y|x)$ and the factor $F(N,p(y|x))$ at 86% of the first training epoch. Both model confidence and factor are evaluated on the unseen data grouped by difficulty level from the MATH dataset. Model confidence decreases with increasing difficulty, whereas the regularization factor increases with problem difficulty. As a result, DCO effectively regularizes contribution from easy examples. This regularization mitigates the potential detrimental effects of overconfidence from easy examples as discussed in Section 4.4. We use $N=256$ corresponding to pass@256 for the plots.

### v3

@@ C.1 Easy data drives overconfidence
C.3 Addition results with Llama-3-70B-base
@@
Figure 8: DCO improves on CE for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier with Llama-3-70B-base. Same as Figure 2 (a) but with Llama-3-70B-base model. We fine-tune Llama-3-70B base models on the MATH dataset to produce direct answers. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. We plot pass@N test coverage as a function of $N$, with each curve (solid red or faint blue-green) corresponding to one fine-tuned model. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
C.4 Out-of-distribution evaluation
@@
Figure 10: DCO improves on CE for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier on out-of-distribution test set. Same as Figure 2 (a) but evaluated on AIME24. We fine-tune Llama-3-8B base models on the MATH dataset to produce direct answers and evaluate on AIME24. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. We plot pass@N test coverage as a function of $N$, with each curve (solid red or faint blue-green) corresponding to one fine-tuned model. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
C.5 Theorem proving
In Table 6, we show additional results for model performance under the DCOSTEP objective. We find that the optimal $N_{\text{eff}}$ grows with increasing passes, agreeing with results in Sections 5.2 and 5.4. We also conduct expert iteration [48, 49] on Mathlib with theorems that do not have proof traces in the training set. We use pass@1k to prove those theorems. We find that our algorithm achieves a stronger improvement over the baseline for pass@4k after the 1st iteration. This improvement might result from the fact that the models can prove more easy theorems where the model has a higher confidence. As a result, we believe our method will perform better with expert iteration.
@@ C.5 Theorem proving
C.6 Plot of the upper bound and the lower bound
@@
C.7 Empirical evaluation of the approximately well calibrated assumption
@@
C.8 The data dependency of confidence and factor $F$
Figure 13: Model is more confident on easy problems; DCO improves test-time scaling by regularizing the easy examples. We fine-tune a Llama-3-8B-base model on the MATH dataset with CE loss and plot the model confidence $p(y|x)$ and the factor $F(N,p(y|x))$ at 86% of the first training epoch. Both model confidence and factor are evaluated on the unseen data grouped by difficulty level from the MATH dataset. Model confidence decreases with increasing difficulty, whereas the regularization factor increases with problem difficulty. As a result, DCO effectively regularizes contribution from easy examples. This regularization mitigates the potential detrimental effects of overconfidence from easy examples as discussed in Section C.1. We use $N=256$ corresponding to pass@256 for the plots.